OpenStudy (anonymous):

Write the equation of the hyperbola whose vertices are at (0,±4) and whose foci are at (0,±11)

OpenStudy (amistre64):

trying for asecond opinion; iwould too ;)

OpenStudy (anonymous):

\[x ^{2}/a -y ^{2}/b =1\]a=b=distance from vertices to center. Difference from center is 4. So your a and b should be 4.

OpenStudy (anonymous):

the transverse axis is vertical.

OpenStudy (anonymous):

The transverse axis is vertical. Thus the standard form of the hyperbola: y^2/a^2-x^2/b^2=1

OpenStudy (anonymous):

Oh yeah, your be is 4, since 4^2 is 16

OpenStudy (anonymous):

b

OpenStudy (amistre64):

11^2 - 4^2 = b^2 right?

OpenStudy (amistre64):

b^2 = 105 then

OpenStudy (amistre64):

c = foci length from center right?

OpenStudy (amistre64):

up 11 and down 11; which to me means the hypotenuse of the a-b triangle

OpenStudy (amistre64):

the transverse in a hyper is got a smaller number under it as well

OpenStudy (anonymous):

hi amistre64 i dont mean to be a bother but I really needed your help for this one question!<3

OpenStudy (amistre64):

the graphics one?

OpenStudy (anonymous):

yes!

OpenStudy (amistre64):

ill try it...

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

c = foci length from center right

OpenStudy (amistre64):

and foci is 11 here; and transverse is 4^2 = 16; so b^2 has to be greater than or equal to 16 at least

OpenStudy (anonymous):

ah, i had a vision problem.. yeah it is 11. i am sorry.

OpenStudy (anonymous):

\[c=11, a=4\] \[b ^{2}=121-16\] \[b ^{2}=105\] So, the equation would be y^2/16 - x^2/(105) =1